OpenStudy (anonymous):

Find a point on the curve y= x^2 that is closest to the point (18, 0).?

OpenStudy (mathmale):

1. I'd suggest graphing this parabola and the given point. 2. Draw (by eye) the shortest line segment connecting (18,0) and the graph of y=x^2. 3. Represent the point at which this line intersects the curve by (x,x^2). 4. Write out an expression for the distance between (18,0) and (x, x^2) (use the distance formula). 5. Differentiate this specific distance formula. 6. Set the result = to 0 7. solve for x, noting that x must be (+) because of its location on the graph of y=x^2. 8. If desired, find the second derivative of this distance and find its sign at this x value. That will tell us whether or not our x-value results in a min. (as desired).

OpenStudy (mathmale):

Having gone through all but the last step, I obtained the equation x-18+2x^3=0, which is easily solved for x. Then the point on the parabola y=x^2 closest to (18,0) is written in terms of that x: (x, x^2).